OpenStudy (anonymous):

Determine whether the following function is even, odd or neither: (Justify your answer in either work or words) f(x) = 2x3 – 3x

OpenStudy (mr.math):

An even function is a function that satisfies \(f(-x)=f(x)\). An odd function is a function that satisfies \(f(-x)=-f(x)\).

OpenStudy (mr.math):

\[f(-x)=2(-x)^3-3(-x)=-2x^3+3x=-(2x^3-3x).\]

OpenStudy (mr.math):

Based on the definitions I gave, what can you conclude?

OpenStudy (mr.math):

Try to compare f(-x) to f(x)!

OpenStudy (mr.math):

Can't you see that \(f(-x)=-f(x)\)?

OpenStudy (anonymous):

so it would be negative, correct?

OpenStudy (mr.math):

What would be negative?

OpenStudy (anonymous):

I'm sorry, I mean positive! The function would be positive?

OpenStudy (mr.math):

Being even and odd is not related to being positive or negative, it doesn't really matter. All you need is to use the definition. For the function you gave, we found that f(-x)=-f(x), and this what an odd function is, thus f(x) is odd.